8. (10 pts.) a) Sketch the region of integration. b) Change the order of integration. [Do not evaluate] SS* 127y dy dx
Answer #1

Solt Given S 12 Wy dydx 715-2 2x44 Y=2x+4 9) (0,4) x=0) x (-2,0) b) 4 - x² = 4 4-y=x² x=4-4 - 9-4 = 2x y - 4 X 2 y-4 2 12 Sy